Q. Lots of 40 components each are called unacceptable if they contain as many as 3 defectives or more. The procedure for sampling the lot is to select 5 components at random and to reject the lot if a defective is found. What is the probability that exactly 1 defective is found in the sample if there are 3 defectives in the entire lot?


1
Expert's answer
2021-05-07T09:23:26-0400

Let X be the number of defective items.


Then using the hypergeometric distribution with "n=5,N=40 \text{ and } k=3"

X~HG(40,5,3)

"P(X=1)=\dfrac{^3C_1\times ^{40-3}C_{5-1}}{^{40}C_5}=\dfrac{^3C_1\times ^{37}C_4}{^{40}C_5}=0.3011"